# Shock waves in nonlinear transmission lines

### v5

@@ VI Half-nonlinear transmission line
For weak shocks, when $|Q_{1}-Q_{2}|$ is small, we can approximated $V(Q)$ for $Q$ between $Q_{1}$ and $Q_{2}$ as
@@
where $\beta$ and $C_{s}$ are constants. We also took $V(Q)$ in the last term in the l.h.s. of (49) in the linear approximation. In this cases, (49) becomes
@@ VII The ODE which doesn’t contain explicitly the independent variable
We would be looking for the solutions of (55) and (60) in terms of elementary functions. Following the well known in mathematics principle, stating that the more general the problem is, the easier it is to solve it, let us change gears and instead of (60) consider the generalized damped Helmholtz-Duffing equation kogan4
@@ VIII Back to the transmission line
Now let us return to the particular cases of Eq. (49) (with the boundary conditions (39a)) considered in Section VI. For (60) using the results of the previous Section we claim that when the boundary conditions satisfy the relation
@@
Note that neither (78) goes to (81) nor (79) goes to (82) when $\gamma\to 0$. Small change in the equation leads to big change of the condition for the solution to be expressed in elementary functions.
We used previously the linear approximation for $V(Q)$ in the last term in the l.h.s. of (49). Strictly speaking, since we are considering nonlinear $V(Q)$ in the r.h.s. of (49) more consistent would be to treat the same way the l.h.s. Equations. (76), (77) allow us to go one step in this direction, that is to consider $V(Q)$ in the last term in the l.h.s. of (49) in quadratic approximation (for cubic nonlinearity of $V(Q)$). The modification of Eqs. (78), (79) can be easily obtained from Eqs. (76), (77) for $n=1$.
@@ Appendix A Shock profile for the fully nonlinear transmission line
If we manage to solve (86), (3) we can substitute the solution into one of the equations (38) and integrate the resulting equation in quadratures to obtain the profile of the shock wave. It is reassuring to realize that the conditions for the integrability in elementary functions of Eqs. (55) and (60) describing weak shocks in half-nonlinear transmission line (i.e. Eqs. (78) and (81)) are actually the (nonlinear) conditions on the product $R_{L}R_{C}$. Further analysis of (86) we postpone until later.

### v6

@@ VI Half-nonlinear transmission line
Equation (49) can be integrated numerically, however in this paper we would be looking for analytical solutions. Trivially, for weak shocks, when $|Q_{1}-Q_{2}|$ is small, we can discard the term with the second derivative in the l.h.s. of the equation and it can be integrated in quadratures. However we’ll see that (49) can be integrated analytically (and even in terms of elementary functions) also in the case when both terms in the l.h.s. of the equation are of the same order of magnitude, though under additional assumptions.
Let us approximate $V(Q)$ for $Q$ between $Q_{1}$ and $Q_{2}$ as
@@
where $\beta$ and $C_{s}$ are constants. Let us also take $V(Q)$ in the last term in the l.h.s. of (49) in the linear approximation. In this cases, (49) becomes
@@ VII The ODE which doesn’t contain explicitly the independent variable
Following the well known in mathematics principle, stating that the more general the problem is, the easier it is to solve it, let us change gears and instead of (60) consider the generalized damped Helmholtz-Duffing equation kogan4
@@ VIII Back to the transmission line
Now let us return to the transmission line. For (60), using the results of the previous Section, we claim that when the parameters of the equation satisfy the relation
@@
Note that (79) doesn’t goes to (82) when $\gamma\to 0$. More specifically, when $\gamma\to 0$, the solution (79) goes to the weak shock solution. On the other hand, the solution (82) corresponds to the case when both terms in (55) are of the same order of magnitude.
We used previously the linear approximation for $V(Q)$ in the last term in the l.h.s. of (49). Strictly speaking, since we are considering nonlinear $V(Q)$ in the r.h.s. of (49) more consistent would be to treat the same way the l.h.s. Equations. (76), (77) allow us to go one step in this direction, that is to consider $V(Q)$ in the last term in the l.h.s. of (49) in quadratic approximation (for cubic nonlinearity of $V(Q)$). As the result, Eqs. (78) and (79) slightly change and Eq. (79) doesn’t change at all.
@@ Appendix A Shock profile for the fully nonlinear transmission line
If we manage to solve (86), (3) we can substitute the solution into one of the equations (38) and integrate the resulting equation in quadratures to obtain the profile of the shock wave. It is reassuring to realize that the conditions for the integrability in elementary functions of Eqs. (55) and (60) describing the shocks in half-nonlinear transmission line (i.e. Eqs. (78) and (81)) are actually the (nonlinear) conditions on the product $R_{L}R_{C}$. Further analysis of (86) we postpone until later.